Hello, My name is Sheila, and I would really like your help in solving this problem, and how to set it up.
The problem is a follow: How many one-to-one correspondence are there between the sets { x,y,z,u,v ] and {1,2,3,4,5 } if in each correspondence:
A. x must correspond to 5
B. x must correspond to 5 and y to 1
C. x,y,z must correspond to odd numbers

Answer by Edwin McCravy(20056)

 { x,y,z,u,v ] and {1,2,3,4,5 }
--------------------------------------------
A. x must correspond to 5

{(x,5),(y,_),(z,_),(u,_),(v,_)}

We can choose 
y's 2nd coordinate 4 ways,
z's 2nd coordinate 3 ways,
u's 2nd coordinate 2 ways,
v's 2nd coordinate 1 way.

Answer: 4·3·2·1 = 4! = 24 ways
--------------------------------------------
B. x must correspond to 5 and y to 1

{(x,5),(y,1),(z,_),(u,_),(v,_)}

We can choose
z's 2nd coordinate 3 ways,
u's 2nd coordinate 2 ways,
v's 2nd coordinate 1 way.

Answer: 3·2·1 = 3! = 6 ways
--------------------------------------------
C. x,y,z must correspond to odd numbers

{(x,_),(y,_),(z,_),(u,_),(v,_)}

We can choose 
x's 2nd coordinate 3 ways, (1, 3 or 5)
y's 2nd coordinate 2 ways,
z's 2nd coordinate 1 way.
u's 2nd coordinate 2 ways, (2 or 4)
v's 2nd coordinate 1 way.

Answer: 3·2·1·2·1 = 3!2! = 6·2 = 12 ways

Edwin